1.In 6 years time I will be twice as old as I was 8 years ago. Determine my age.
2.Bob is twice as old as Alice at the moment. In 4 years time their total age will be 71. What is Alice’s age now?
3.A Triple Jump consists of a hop, step and jump. The length of Keith’s step was three-quarters of the length of his hop and the length of his jump was half the length of his step. If the total length of Keith’s triple jump was 17m, what was the length of his hop, in metres?

Answer by ikleyn(52788)

.
1.In 6 years time I will be twice as old as I was 8 years ago. Determine my age.
2.Bob is twice as old as Alice at the moment. In 4 years time their total age will be 71. What is Alice’s age now?
3.A Triple Jump consists of a hop, step and jump. The length of Keith’s step was three-quarters of the length of his hop and the length of his jump was half the length of his step. If the total length of Keith’s triple jump was 17m, what was the length of his hop, in metres?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

I will solve #1 only.

Let x be your age now.

Then the condition says that

x + 6 = 2(x-8).

Simplify and solve:

x + 6 = 2x - 16  --->  6 + 16 = 2x - x  --->  22 = x.

Answer. You are 22 years old now.

Check.  x + 6 = 22 + 6 = 28.
               x - 8 = 22 - 8 = 14.    28 = 2*14.  Correct!